how do you solve this 150/2.5 = x/6?
Answer by vincygrrl22(7)

150/2.5=x/6
simplify both sides of the equation:
375=x/6
Multiply both sides of the equation by 6:
60*6 = (x/6)6
360=x
OR you can cross multipy:
150*6=2.5x
then divide both sides by 2.5
900/2.5=(2.5x)/2.5=
360=x
X=360